OpenStudy (anonymous):

NEED HELP.... Calculate the integration of sin^3(X)cos^6(X)dx....

OpenStudy (cruffo):

\[\int \sin^3(x)\; \cos^6(x) \;dx\] Looks like chain rule may be involved. Have you tried u-substitution?

OpenStudy (anonymous):

i did the u subsitution...but im stuck on it .. no idea where i went wrong

OpenStudy (cruffo):

maybe we need to reduce the power on cos a little. Some trig identity???

OpenStudy (anonymous):

wouldnt you reduce the sin since its an odd exponent..

OpenStudy (anonymous):

then use cos^2x+sin^2x=1....

OpenStudy (cruffo):

yah, \[\sin^3(x) = \sin(x)\sin^2(x) = \sin(x) (1-cos^2(x))\] so at least one part of the integral can be dealt with...

OpenStudy (anonymous):

i got = (1-u^2)(u^6)(-du).....

OpenStudy (anonymous):

that was for the subsitution before integrating... \

OpenStudy (anonymous):

u = cosx....du=-sinx

OpenStudy (cruffo):

ok...

OpenStudy (anonymous):

did some distributing before i integrateded... and got u^12-u^6 du...

OpenStudy (cruffo):

??? I got \[\int (u^6 - u^8) du\]

OpenStudy (anonymous):

ohhhhh you add exponents ... thats right... forgot

OpenStudy (cruffo):

:)

OpenStudy (anonymous):

that was my flaw.... thank you

OpenStudy (cruffo):

cool